OpenStudy (anonymous):

Your piggy bank has a total of 18 coins in it; some are dimes and some are nickels. If you have a total of $1.40, how many nickels do you have?

OpenStudy (klimenkov):

Let there be \(x\) nickels and \(y\) dimes. So we have 2 equations: \[x+y=18\]\[5x+10y=140\]All you need is to solve this sytem.

OpenStudy (anonymous):

Could you tell me what's the worth of nickels and dimes?

OpenStudy (anonymous):

\[\huge \color{green}{\textbf{Welcome To Openstudy..}}\]

OpenStudy (anonymous):

See, you are given with : x + y = 18 Evaluate y from here by subtracting x from both the sides, y = 18 - x Now: 5x + 10 y = 140 5x + 10(18 - x) = 140 5x - 10x + 180 = 140 5x = 40 x = 8 So, he has 8 nickels..